# Nuclear Cold QCD: Review and Future Strategy

### v1

@@ V.2.2 Breakup by Comovers
The absorption models just discussed do not generally distinguish between $J/\psi$ and $\psi^{\prime}$ suppression at forward $x_{F}$ because, at this point, the final quarkonium state is fully formed outside the nucleus. However, absorption by comovers, whether these comovers are characterized as hadrons or partons [23], does do so because it assumes that the comovers interact with formed charmonium states, consistent with larger interaction cross sections for the $\psi^{\prime}$, especially at energies where the nucleon absorption cross section is assumed to be negligible [425] but comover suppression is active [427].
In the comover model [428, 429, 430, 431, 432, 433], the suppression arises from scattering of the nascent charmonium with produced particles that travel along the path of the $c\bar{c}$ [428, 434]. The comover survival probability also depends on the time during which the charmonium state can interact with the comovers. The result depends on the ratio $\tau_{f}/\tau_{0}$, the final over the initial time. Because the proper time is inversely proportional to the density, $\tau_{f}/\tau_{0}=\rho^{\rm co}(\vec{b},z)/\rho_{pp}$, where the final time at which the interactions cease is when the comover density has decreased to the particle density in $p+p$ collisions at the same energy. Thus,
@@
The values of $\sigma_{\rm co}$ were fixed [429] from fits to fixed-target data to be $\sigma_{\rm co}^{J/\psi}=0.65$ mb and $\sigma_{\rm co}^{\psi^{\prime}}=6$ mb. These same values have also been used to describe data from RHIC [435] and the LHC [436].
@@ V.3 Observables
Several observables can help isolate collisional dissociation and absorption effects, if they exist. These are also particularly useful for distinguishing dissociation and absorption from other competing suppression mechanisms such as energy loss or nPDF modifications.
@@
  • In dissociation calculations, once the strength of strength of the heavy quarks in matter is set, [437] $\mathcal{R}$ can be predicted for individual quarkonium states based on the width of the quarkonium wavefunction. More weakly bound quarkonium states experience more suppression than tightly bound ones [419, 420], following a hierarchy that can be quantitatively tested, see the right panel of Fig. 17.
  • The $\chi_{c}$ is primarily produced as a color singlet, while the $J/\psi$ and $\psi^{\prime}$(2S) predominantly originate from color octets. As a result, their final-state interactions in nuclei are expected to differ significantly. FCEL and nPDF effects are effectively eliminated in $\mathcal{R}$ in Eq. (16).
  • Finally, hadron production in semi-inclusive DIS $e+A$ collisions can be considered. The ratio between hadron yields associated with electrons scattered off nuclei and protons, can isolate absorption effects if final-state energy loss is small for high energy hadrons.
@@ VI Other nonperturbative effects
In this section, we focus on two non-perturbative initial-state effects. The first concerns the role of the impact parameter in leading-twist nuclear shadowing in the nPDF framework in $h+A$ and $e+A$ collisions. So far, the spatial dependence—i.e., the location within the nucleus where the probe interacts—has largely been neglected in standard calculations. The second non-perturbative effect involves intrinsic charm, whose existence has recently received renewed support from LHCb measurements [438].
@@ VI.1 Impact Parameter Dependent Shadowing
The nuclear modifications of the parton densities are derived from nuclear deep-inelastic scattering data and minimum-bias nuclear collisions, as in EPPS21 [4]. However, given that the parton distributions are modified in the presence of the nuclear medium, it is also reasonable to assume that the effect may further depend on where in the nucleus the parton originated, from a nucleon closer to the nuclear surface or embedded deeply in its volume. No detailed studies of parton distributions in the nucleus as a function of collision centralities exist so far. The first experimental study of spatial dependence of parton distributions relied on dark tracks in emulsion to tag more central collisions [439]. Studies of the impact parameter dependence of shadowing were first proposed in Ref. [440].
@@
where $s=\sqrt{b^{2}+z^{2}}$, $b$ is the transverse impact parameter, $z$ the longitudinal position, and $k_{T}$ the transverse momentum. If the parton densities in the nucleon and nucleus are the same, then $S_{i}=1$.
@@
where $R_{A}$ is the nuclear radius, $d$ the surface thickness, and $\omega$ allows for central density variation.
Two shadowing parameterizations were proposed in Ref. [440], giving stronger shadowing at low $b$ (central collisions) and weaker at high $b$ (peripheral collisions), normalized so that
@@
where $N_{\rm WS}$ ensures proper normalization. This form has weak $s$-dependence until near the surface.
Alternatively, if shadowing depends on the nuclear thickness $T_{A}(b)$, the $b$-dependence is stronger. In both approaches, $S_{i}(\vec{b},z)<S_{i}$ in the shadowing and EMC regions for small $b$, and $S_{i}(\vec{b},z)>S_{i}$ in the antishadowing region. If no $b$-dependence is assumed, the same shadowing is applied to all $b$, which is unrealistic. As shown in Ref. [441], a model like $S_{i}^{\rm WS}$ enhances shadowing at $b=0$ by a few percent, halves it around $b\sim 1.5R_{A}$, and eliminates it for $b>2R_{A}$. A stronger $b$-dependence would both increase central shadowing and localize it more effectively.
In Ref. [442], these parameterizations were found insufficient to describe the PHENIX d+Au data [443]. Two additional models were tested. EPS09s [444], similar to the previous form but using powers of $T_{A}^{n}(b)$ up to $n=4$ with $A$-independent coefficients $a(n)$,
@@
which provided the best fit with $R=2.4$ fm and $d=0.12$ fm, suggesting a well-localized shadowing zone at the center, like a hot spot [442].
Later ALICE measurements indicated a weaker $b$-dependence, implying that at higher energies, shadowing is more spatially extended across the nucleus rather than localized [445]. However, note that the $b$-dependent analysis in Ref. [445] does not account for other cold nuclear matter effects.
@@ VI.2 Intrinsic charm
The basic five-particle proton state produces leading charm hadrons, carrying valence quarks of the parent hadron, such as the $\overline{D}$ in the proton and $D^{-}$ from a similar four-particle state in the $\pi^{-}$, resulting in asymmetries between $D^{-}$ and $D^{+}$ in $\pi^{-}+p$ interactions, particularly at high $x_{F}$ or rapidity [453, 454, 455, 451]. Intrinsic states with additional light or heavy $q\overline{q}$ pairs have been studied [456, 457], including tetraquark production [458].
@@
The resolving factor $Q^{2}/(4\widehat{m}_{c}^{2})$, arising from a soft interaction that breaks the coherence of the state [452], was introduced to calculate the intrinsic charm cross section in Ref. [464]. Here $Q^{2}=0.1$ GeV2 is assumed [465]. The inelastic cross section, $\sigma_{pN}^{\rm in}=30$ mb, is appropriate for fixed-target interactions and can be used as a conservative estimate at higher energies. The cross sections for higher states can be scaled appropriately by the masses; see the discussion in Ref. [458].
Intrinsic heavy quark production differs from standard perturbatively produced heavy quarks in the treatment of the initial state. As previously noted, the intrinsic production occurs in forward kinematic regions since the constituent quarks of the final-state hadron all emerge from the projectile hadron itself. In contrast, in perturbative production of heavy quarks, e.g., via $gg\rightarrow Q\overline{Q}X$, one gluon arises from the projectile and the other from the target, and the $Q$ and $\overline{Q}$ are produced in the center of the collision with no significant distinction in the kinematics of the final-state $Q$ and $\overline{Q}$.
@@
Here $N$ denotes either a $p$ or $A$ target, with $A=1$ in the $A$-dependent coefficients for $p+p$ collisions. The exponents $\alpha$ and $\beta$ encapsulate the different nuclear dependencies of the two components. In general, kinematic effects due to cold nuclear matter (CNM), such as the modification of the parton densities in the nucleus, energy loss, and momentum broadening, can be incorporated into the perturbative QCD cross section, leaving only final-state effects such as nuclear absorption or comover suppression to be included in $\alpha$.
@@
Intrinsic charm distributions have been included in global analyses of the parton densities [462, 469, 470, 471, 472, 473] with contradictory results. The NNPDF collaboration established the existence of intrinsic charm to the $3\sigma$ level in their analysis [474], consistent with both the LHCb $Z+$charm results [448] and the previous EMC charm quark measurements [475]. However, other work has called the NNPDF results into question based on how the LHCb $Z+c$-jet data were analyzed [476].
@@ VI.3 Observables
  • The impact parameter dependence of the nPDFs has not been studied in any systematic way so far. Most of the focus has so far been on studies of quarkonium production in $h+A$ collisions as a function of centrality-related variables. The PHENIX d+Au data [443] studied in Ref. [442] suggested a stronger effect than ALICE charmonium centrality-dependent measurements in $p+{\rm Pb}$ collisions [479, 480]. The CMS Collaboration [481] has studied $\Upsilon$ production as a function of event activity. Differences in the RHIC and LHC results could arise from changes in the gluon density of the nucleus as higher energies probe lower $x$ regions and more central collisions could herald a transition from a more dilute system with small hot spots as suggested in Ref. [442] to a more fully saturated system of gluons.
@@ VII Future Experimental Directions
The systematic study of Drell–Yan production and other hard probes such as heavy flavor and quarkonia in $p+A$ collisions as a function of $A$, $x_{F}$, and $p_{T}$, is crucial for disentangling the mechanisms of CNM energy loss and the modification of nuclear PDFs, as welll as other CNM effects. Ongoing and future experiments, such as SeaQuest/SpinQuest at Fermilab [26] and AMBER at CERN [478], will provide increasingly precise data to test theoretical models and improve our understanding of quark propagation and energy loss in cold nuclear matter and separate these contributions from nPDF effects. The proposed Forward Physics Facility at CERN [483] could also contribute significantly to the understanding of proton and nuclear structure, coherent power corrections, and quantum electrodynamics (QED) corrections.
@@ VII.1 Fixed-Target Experiments
At Fermilab, the SeaQuest experiment (E906) [484], which concluded data taking in 2017, with ongoing analyses, used a 120 GeV proton beam to study Drell–Yan production with several nuclear targets (carbon, iron, tungsten) along with reference targets (liquid hydrogen, deuterium). By mapping the Drell–Yan cross section across a range of nuclear masses, SeaQuest can separate partonic energy loss from other CNM effects, such as nuclear shadowing, while refining nPDFs in the medium-to-high $x$ region. Its successor, SpinQuest (E1039) [485], collects data since 2019. SpinQuest focuses on spin-dependent phenomena with polarized targets, exploring the Sivers function, which provides insights into the orbital angular momentum of sea quarks. Future phases may include nuclear targets to enhance CNM physics measurements.
At CERN, AMBER (NA66) [478] will extend the legacy of COMPASS [486] by exploring hadronic structure and nuclear effects with high‐intensity pion, kaon, and proton beams from 60 to 190 GeV. AMBER’s Drell–Yan and charmonium measurements on nuclear targets will place high‐precision constraints on initial‐state energy loss and distinguish between coherent and incoherent mechanisms. Its broad kinematic coverage will also refine our understanding of nPDFs.
The proposed NA60+ experiment at CERN [477] (an extension of NA60), aims to measure dimuon production with improved precision with proton beams of 40, 80 and 120 GeV. Their high precision measurements of quarkonium and Drell–Yan at these low energies will offer a more comprehensive energy dependence of partonic interactions in nuclear medium. The NA61/SHINE experiment [487], taking data since 2009, although primarily focused on hadron yields, could provide additional insights into nuclear PDFs and energy loss mechanisms in dedicated runs enabled by the post-2020 upgrade proposal [488].
At the LHC, the SMOG2 upgrade at LHCb [489, 490] provides a unique approach to fixed‐target physics at TeV energies. By injecting low‐pressure gas into the beam pipe, SMOG2 enables $p+A$ and ${\rm Pb}+A$ collisions at forward rapidities, providing a unique environment for probing CNM effects. Although Drell–Yan measurements may be limited by relatively low integrated luminosities, they would complement quarkonium and open heavy flavor studies, especially for energy-loss dynamics at high fixed–target energies [491, 492]. Further measurements of asymmetries between leading and nonleading $D$ mesons with SMOG2 could help separate possible contributions from intrinsic charm with other hadronization mechanisms [468].

### v2

@@ V.2.2 Breakup by Comovers
The absorption models just discussed do not generally distinguish between $J/\psi$ and $\psi$(2S) suppression at forward $x_{F}$ because, at this point, the final quarkonium state is fully formed well outside the nucleus. However, absorption by comovers, whether these comovers are characterized as hadrons or partons [23], does do so because it assumes that the comovers interact with formed charmonium states, consistent with larger interaction cross sections for the $\psi$(2S), especially at energies where the nucleon absorption cross section is assumed to be negligible [427] but comover suppression is active [429].
In the comover model [430, 431, 432, 433, 434, 435], the suppression arises from scattering of the nascent charmonium with produced particles that travel along the path of the $c\bar{c}$ [430, 436]. The comover survival probability also depends on the time during which the charmonium state can interact with the comovers. The result depends on the ratio $\tau_{f}/\tau_{0}$, the final over the initial time, effectively the lifetime of the system. Because the proper time is inversely proportional to the density, $\tau_{f}/\tau_{0}=\rho^{\rm co}(\vec{b},z)/\rho_{pp}$, where the final time at which the interactions cease is when the comover density has decreased to the particle density in $p+p$ collisions at the same energy. Thus,
@@
The values of $\sigma_{\rm co}$ were fixed [431] from fits to fixed-target data to be $\sigma_{\rm co}^{J/\psi}=0.65$ mb and $\sigma_{\rm co}^{\psi{\rm(2S)}}=6$ mb. These same values have also been used to describe data from RHIC [437] and the LHC [438]. Comover suppression has also been applied to excited states of bottomonium at the LHC [439]. Comover suppression can explain differences in suppression patterns between excited quarkonium states in $p+A$ collisions, particularly where absorption effects should be small. Since the comover effects depend on the density of the produced particles in the collision, this mechanism results in different dependencies on charged particle multiplicity or impact parameter.
@@ V.3 Observables
Several observables can help isolate collisional dissociation, absorption, and comover effects, if they exist. These are also particularly useful for distinguishing these effects from other competing suppression mechanisms such as energy loss or nPDF modifications which are independent of whether the quarkonia are produced in an excited state or the ground state.
@@
  • In higher energy collisions, where $t_{\rm form}\gg L$ and absorption is likely negligible, measurements of $\mathcal{R}$ as a function of the density of produced particles could help isolate comover effects. In addition, since comover suppression is expected to apply only to quarkonium production, a comparison between $\mathcal{R}$ in Eq. (16) and double ratios of open heavy flavor mesons could be helpful.
  • In calculations of collisional dissociation, once the binding energy of the heavy quarks in matter is fixed [440], $\mathcal{R}$ can be predicted for individual quarkonium states based on the width of the quarkonium wavefunction. More weakly bound quarkonium states experience more suppression than tightly bound ones [421, 422], following a hierarchy that can be quantitatively tested, see the bottom panel of Fig. 17.
  • The $\chi_{c}$ is primarily produced as a color singlet, while the $J/\psi$ and $\psi$(2S) predominantly originate from color octets. As a result, their final-state interactions in nuclei are expected to differ significantly. FCEL and nPDF effects are effectively eliminated in $\mathcal{R}$ in Eq. (16).
  • Finally, the hadron production ratio between $e+A$ and $e+p$ in semi-inclusive DIS can isolate absorption effects if final-state energy loss is small for high energy hadrons.
@@ VI Other nonperturbative effects
In this section, we focus on two nonperturbative initial-state effects. The first concerns the role of the impact parameter in leading-twist nuclear shadowing in the nPDF framework in $h+A$ and $e+A$ collisions. So far, the spatial dependence, the location within the nucleus where the probe interacts, has largely been neglected in standard calculations. The second nonperturbative effect is intrinsic charm, the existence of which has recently received renewed support from LHCb measurements [441].
@@ VI.1 Impact Parameter Dependent Shadowing
The nuclear modifications of the parton densities are derived from nuclear DIS data and minimum-bias $h+A$ collisions, as in EPPS21 [4]. However, given that the parton distributions are modified in the presence of the nuclear medium, it is also reasonable to assume that the effect may further depend on where in the nucleus the parton originated, from a nucleon closer to the nuclear surface or embedded deeply in its volume. No detailed studies of parton distributions in the nucleus as a function of collision centralities exist so far. The first experimental study of spatial dependence of parton distributions relied on dark tracks in emulsion to tag more central collisions [442]. Studies of the impact parameter dependence of shadowing were first proposed in Ref. [443].
@@
where $s=\sqrt{b^{2}+z^{2}}$, $b$ is the transverse impact parameter, $z$ the longitudinal position, and $k_{T}$ the transverse momentum. If the parton densities in the nucleon and nucleus are the same, then $S_{i}=1$. Then the integration over the nuclear volume gives $F_{i}^{A}(x,Q^{2},k_{T})=Af_{i}^{p}(x,Q^{2},k_{T})$.
@@
where $R_{A}$ is the nuclear radius, $d$ the surface thickness, $\omega$ allows for central density variation, and $\rho_{0}$ is the density at the center of the nucleus.
Two shadowing parameterizations were proposed in Ref. [443], both giving stronger shadowing at low $b$ (central collisions) and weaker at high $b$ (peripheral collisions), normalized so that
@@
where $N_{\rm WS}$ ensures proper normalization. This form has weak spatial dependence until near the surface.
Alternatively, if shadowing depends on the nuclear thickness $T_{A}(b)$, the $b$ dependence is stronger. In both approaches, $S_{i}(\vec{b},z)<S_{i}$ in the shadowing and EMC regions for small $b$, and $S_{i}(\vec{b},z)>S_{i}$ in the antishadowing region. If no $b$ dependence is assumed, the same shadowing is applied to all $b$, which is unrealistic. As shown in Ref. [444], a model like $S_{i}^{\rm WS}$ enhances shadowing at $b=0$ by a few percent, halves it around $b\sim 1.5R_{A}$, and eliminates it for $b>2R_{A}$. A stronger $b$ dependence would both increase central shadowing and localize it more effectively.
In Ref. [445], these parameterizations were found insufficient to describe the PHENIX d+Au data [446]. Two additional models were tested. EPS09s [447], similar to the $T_{A}$-dependent form but using powers of $T_{A}^{n}(b)$ up to $n=4$ with $A$-independent coefficients $a(n)$,
@@
which provided the best fit with $R=2.4$ fm and $d=0.12$ fm, suggesting that shadowing was localized at the center of the nucleus, like a hot spot [445].
Later ALICE measurements indicated a weaker $b$ dependence, implying that at higher energies, shadowing is more uniformly distributed throughout the nuclear volume rather than localized at the center [448]. However, the $b$-dependent analysis in Ref. [448] does not account for other cold nuclear matter effects.
@@ VI.2 Intrinsic charm
The basic five-particle proton state produces leading charm hadrons, carrying valence quarks of the parent hadron, such as the $\overline{D}$ in the proton and $D^{-}$ from a similar four-particle state in the $\pi^{-}$, resulting in asymmetries between $D^{-}$ and $D^{+}$ in $\pi^{-}+p$ interactions, particularly at high $x_{F}$ or rapidity [456, 457, 458, 454]. Intrinsic states with additional light or heavy $q\overline{q}$ pairs have been studied [459, 460], including tetraquark production [461, 462].
@@
The resolving factor $\mu^{2}/(4\widehat{m}_{c}^{2})$, arising from a soft interaction that breaks the coherence of the state [455], was introduced to calculate the intrinsic charm cross section in Ref. [468]. Here $\mu^{2}=0.1$ GeV2 is assumed [469]. The inelastic cross section, $\sigma_{pN}^{\rm in}=30$ mb, is appropriate for fixed-target interactions and can be used as a conservative estimate at higher energies. The cross sections for higher states can be scaled appropriately by the masses; see the discussion in Ref. [461].
Intrinsic heavy quark production differs from standard perturbatively produced heavy quarks in the treatment of the initial state. As previously noted, intrinsic production occurs in forward kinematic regions since the constituent quarks of the final-state hadron all emerge from the projectile hadron itself. In contrast, in perturbative production of heavy quarks, e.g., via $gg\rightarrow Q\overline{Q}X$, one gluon arises from the projectile and the other from the target, and the $Q$ and $\overline{Q}$ are produced in the center of the collision with no significant distinction between the kinematics of the final-state $Q$ and $\overline{Q}$.
@@
Here $N$ denotes either a $p$ or $A$ target, with $A=1$ in the $A$-dependent coefficients for $p+p$ collisions. The exponents $\alpha$ and $\beta$ encapsulate the different nuclear dependencies of the two components. In general, CNM kinematic effects, such as nPDFs, energy loss, and momentum broadening, can be incorporated into the perturbative QCD cross section, leaving only final-state effects such as nuclear absorption or comover suppression to be included in $\alpha$.
@@
Intrinsic charm distributions have been included in global analyses of the parton densities [466, 473, 474, 475, 476, 477] with contradictory results. The NNPDF collaboration established the existence of intrinsic charm to the $3\sigma$ level in their analysis [478], consistent with both the LHCb $Z+$charm results [451] and the previous EMC charm quark structure function, $F_{2}^{c}$ measurements [479]. However, other work has called the NNPDF results into question based on how the LHCb $Z+c$-jet data were analyzed [480].
@@ VI.3 Observables
  • The impact parameter dependence of the nPDFs has not been studied in any systematic way so far. Most of the focus has been on studies of quarkonium production in $h+A$ collisions as a function of centrality-related variables. The PHENIX d+Au data [446] studied in Ref. [445] suggested a stronger effect than ALICE charmonium centrality-dependent measurements in $p+{\rm Pb}$ collisions [483, 484]. The CMS collaboration [485] has studied $\Upsilon$ production as a function of event activity. Differences in the RHIC and LHC results could arise from changes in the gluon density of the nucleus as higher energies probe lower $x$ regions and more central collisions could herald a transition from a more dilute system with small hot spots as suggested in Ref. [445] to a more fully saturated system of gluons.
@@ VII Future Experimental Directions
The systematic study of Drell–Yan production and other hard probes such as heavy flavor and quarkonia in $p+A$ collisions as a function of $A$, $x_{F}$, and $p_{T}$, is crucial for disentangling the mechanisms of CNM energy loss and the modification of nuclear PDFs, as well as other CNM effects. Ongoing and future experiments, such as SeaQuest/SpinQuest at Fermilab [26] and AMBER at CERN [482], will provide increasingly precise data to test theoretical models and improve our understanding of quark propagation and energy loss in cold nuclear matter and separate these contributions from nPDF effects. The proposed Forward Physics Facility at CERN [487] could also contribute significantly to the understanding of proton and nuclear structure, coherent power corrections, and quantum electrodynamics (QED) corrections.
@@ VII.1 Fixed-Target Experiments
At Fermilab, the SeaQuest experiment (E906) [488], which concluded data taking in 2017, with ongoing analyses, used a 120 GeV proton beam to study Drell–Yan production with several nuclear targets (carbon, iron, tungsten) along with reference targets (liquid hydrogen, deuterium). By mapping the Drell–Yan cross section across a range of nuclear masses, SeaQuest can separate partonic energy loss from other CNM effects, such as nuclear shadowing, while refining nPDFs in the medium-to-high $x$ region. SeaQuest also studied $J/\psi$ production and could provide insight into intrinsic charm contributions [469]. Its successor, SpinQuest (E1039) [489], collects data since 2019. SpinQuest focuses on spin-dependent phenomena with polarized targets, exploring the Sivers function, which provides insights into the orbital angular momentum of sea quarks. Future phases may include nuclear targets to enhance CNM physics measurements.
At CERN, AMBER (NA66) [482] will extend the legacy of COMPASS [490] by exploring hadronic structure and nuclear effects with high‐intensity pion, kaon, and proton beams from 60 to 190 GeV. AMBER’s Drell–Yan and charmonium measurements on nuclear targets will place high‐precision constraints on initial‐state energy loss and distinguish between coherent and incoherent mechanisms. Its broad kinematic coverage will also refine our understanding of the nPDFs.
The proposed NA60+ experiment at CERN [481] (an extension of NA60), aims to measure dimuon production with improved precision with proton beams of 40, 80 and 120 GeV. Their high precision measurements of quarkonium and Drell–Yan at these low energies will offer a more comprehensive energy dependence of partonic interactions in a nuclear medium. The NA61/SHINE experiment [491], taking data since 2009, although primarily focused on hadron yields, could provide additional insights into nuclear PDFs and energy loss mechanisms in dedicated runs enabled by the post-2020 upgrade proposal [492].
At the LHC, the SMOG2 upgrade at LHCb [493, 494] provides a unique approach to fixed‐target physics at TeV energies. By injecting low‐pressure gas into the beam pipe, SMOG2 enables $p+A$ and ${\rm Pb}+A$ collisions at forward rapidities, providing a unique environment for probing CNM effects. Although Drell–Yan measurements may be limited by relatively low integrated luminosities, they would complement quarkonium and open heavy flavor studies, especially for energy-loss dynamics at high fixed–target energies [495, 496]. Further measurements of asymmetries between leading and nonleading $D$ mesons with SMOG2 could help separate possible intrinsic charm contributions from other hadronization mechanisms [472].